(Px-y)(Py+x)=2P


1
Expert's answer
2021-05-07T09:05:02-0400

(px-y)(x-py)=2p

(pxy-y2)(x-yp)=2yp

(py/x.x2-y2)(1-py/x=2py/x

py/x.x 2-y2=[2py/x]/1-py/x

y2=py/x.x2-[2py/x]/[1-py/x]

then x2=u,y2=v

du=2xdx,dv=2ydy

p=dy/dx=2x/2y.dv/du=u/v.p,p=dv/du

py/x=dv/du=P

substitute

v=Pu-2P/1-P

v=Pu+ f(9)

v=Pu- f(p)

the solution is given by,

v=Cu-2C/1-C,c is arbitrary constant.